OpenStudy (anonymous):

Cell phone coverage is in a circular pattern form the tower. if the equation of this pattern is x^2+y^2=2025. How many miles away from the tower will you have coverage?

OpenStudy (danjs):

x^2 + y^2 = (radius)^2

OpenStudy (danjs):

you want the radius, that is the coverage distance

OpenStudy (anonymous):

Ok thanks I had no idea where to start.

OpenStudy (danjs):

The equation for a circle, is x^2 + y^2 = (radius)^2 Here they just give you 2025

OpenStudy (danjs):

(radius)^2 = 2025

OpenStudy (danjs):

the coverage distance will be the square root of 2025

OpenStudy (anonymous):

Ok thanks again.

OpenStudy (danjs):

x^2 + y^2 = r^2 x^2 + y^2 = 2025 = 45^2

OpenStudy (danjs):

yw